# SlowFast-LLaVA: A strong training-free baseline for video large language models

### v1

@@ 4.3 Main Results
We compare SF-LLaVA against state-of-the-art training-free methods, such as IG-VLM (Kim et al., 2024) and LLoVi (Zhang et al., 2023a), on multiple VideoQA tasks. In addition, we also compare with other Video LLMs that are supervised fine-tuned (SFT) with video datasets, such as Video-LLaVA (Lin et al., 2023) and PLLaVA (Xu et al., 2024). This comparison allows us to illustrate our position within the broader landscape of existing Video LLMs.
Open-Ended VideoQA results are shown in Table 1. SF-LLaVA obtains better performance than state-of-the-art training-free methods on all benchmarks. Specifically, SF-LLaVA outperforms IG-VLM (Kim et al., 2024) by 2.1% and 5.0% on MSRVTT-QA, 5.7% and 1.5% on TGIF-QA, 2.0% and 0.8% on ActivityNet-QA, when using 7B and 34B LLMs, respectively. When even compared to SFT methods, SF-LLaVA achieves on-par results on most benchmarks (i.e., MSVD-QA, MSRVTT-QA, and TGIF-QA), and only the results of PLLaVA (Xu et al., 2024) and LLaVA-NeXT-Video-DPO (Zhang et al., 2024b) are better than ours on ActivityNet-QA.
Multiple Choice VideoQA results are shown in Table 2. SF-LLaVA outperforms other training-free methods that use the comparable LLMs and visual encoders, such as IG-VLM (Kim et al., 2024) on all benchmarks. Specifically, on the challenging EgoSchema dataset, which involves complex long-form temporal reasoning, (Mangalam et al., 2024), SF-LLaVA outperforms IG-VLM by 11.4% and 2.2% when using 7B and 34B LLMs, respectively. This highlights the ability of SF-LLaVA on long-form video understanding. Note that VideoTree (Wang et al., 2024b) is leading the benchmark because it is built upon a proprietary LLM (i.e., GPT-4 (Achiam et al., 2023)) whose performance is much better than the open-sourced LLMs. When compared to SFT methods (Cheng et al., 2024), SF-LLaVA 34B model also achieves better results (+2.5%) on EgoSchema, which confirms the capability of our SlowFast design on long videos.
@@
Text Generation benchmarks are shown in Table 3, where SF-LLaVA still demonstrates some advantages. Specifically, SF-LLaVA-34B outperforms all training-free baselines on average. We observe that SF-LLaVA consistently performs worse than LLaVA-NeXT-Image (Zhang et al., 2024b) on Detail Orientation (DO). This is because LLaVA-NeXT-Image takes as inputs more “high-resolution” frames than ours (i.e., 32 frames with $12\times 12$ tokens v.s. 10 frames with $12\times 24$ tokens), thus is able to capture more spatial information along the video. On the other hand, SF-LLaVA takes advantage of the SlowFast design to cover longer temporal context by using even less visual tokens (i.e., 4608 tokens v.s. 3680 tokens), thus excels in all other tasks, especially on Temporal Understanding (TU). We also observe that SF-LLaVA-34B is superior to most SFT methods (e.g, outperforming Video-LLaMA2 (Cheng et al., 2024) +0.1 score on TU and +0.31 score on CO), but only needs to catch up with LLaVA-NeXT-Video-DPO (Zhang et al., 2024b).
@@ 4.4 Design Choices of SlowFast
We first validate if both the Slow and Fast pathways are necessary, and then continue to experiment for their design choices, respectively. These ablation studies are conducted on ActivityNet-QA (an Open-Ended VideoQA dataset that contains videos of human activities) and EgoSchema (a Multiple Choice VideoQA dataset that requires long-form understanding of egocentric videos).
@@
Can we remove the Slow pathway? We remove the Slow pathway to evaluate its contribution. First, we simply remove the Slow pathway, while keeping the Fast pathway as 50 frames, each with $4\times 4$ tokens. Fig. 3 shows that, on all benchmarks, removing Slow pathway (number of frames equals to 0) will introduce much lower performance than keeping it. Second, we validate if the performance gain is caused by the necessity of the Slow pathway or the increased visual tokens brought by more frames. We test to remove the Slow pathway and gradually increase $N^{\textnormal{fast}}$ from 50 to 100, 125, 150, 175, 200 and 225, to compensate the number of visual tokens. The results are summarized in Table 4. We observe that using larger $N^{\textnormal{fast}}$ generally obtains better results, but the performance quickly saturate when $N^{\textnormal{fast}}$ is larger than 150 (e.g., using 175 and 200 frames get on-par results on both ActivityNet-QA and EgoSchema). We also compare the baselines in Table 4, which use $N^{\textnormal{fast}}=200$, and SF-LLaVA models in Fig. 3 with $N^{\textnormal{slow}}=8$ and $N^{\textnormal{fast}}=50$, since these models use similar number of tokens ($200\times 4\times 4=3200$ v.s.$8\times 12\times 24+50\times 4\times 4=3104$) in total. Results show that SF-LLaVA outperforms this new baseline under all settings (e.g., 54.7% v.s. 49.7% and 46.0% v.s. 37.0% using 7B LLMs on ActivityNet-QA and EgoSchema, respectively). All of above results demonstrate that it is essential to use the Slow pathway in SF-LLaVA.
@@
Table 4: Effect of increasing the number of frames in the Fast pathway each with $4\times 4$ tokens. Symbol “–” means the model inference gets out-of-memory on 80GB GPUs under this setting.
Can we remove the Fast pathway? We validate this by removing the Fast pathway while retaining the Slow pathway (using 10 frames and $12\times 24$ tokens for each frame). Fig. 4 shows that SF-LLaVA with $N^{\textnormal{fast}}=50$ consistently outperforms this baseline on all benchmarks. Similar to the experiments for the Slow pathway, we increase $N^{\textnormal{slow}}$ to ensure SF-LLaVA and this new baseline with only Slow-pathway having a comparable number of input video tokens. Specifically, we increase $N^{\textnormal{slow}}$ to 12 frames, which is the maximum number of frames that the 34B model can afford. SF-LLaVA still outperforms this new baseline on both ActivityNet-QA (56.3% v.s. 54.3% on 7B model and 59.2% v.s. 58.8% on 34B model) and EgoSchema (47.2% v.s. 46.6% on 7B model and 55.8% v.s. 54.6% on 34B model). We observe that the performance gap is more obvious on EgoSchema. This is because EgoSchema (Mangalam et al., 2024) mostly contains long-form videos and answering the questions requires to capture longer context using the Fast pathway.
@@
7B $10\times 24\times 24$ 53.7/3.3 53.2
$10\times 12\times 24$ 54.8/3.3 47.2
$10\times 12\times 12$ 54.3/3.3 39.0
$5\times 24\times 24$ 54.1/3.3 44.4
$5\times 12\times 24$ 53.6/3.3 39.4
$5\times 12\times 12$ 51.2/3.2 36.4
@@
$10\times 12\times 24$ 58.9/3.5 55.8
$10\times 12\times 12$ 58.3/3.5 50.8
$5\times 24\times 24$ 58.0/3.5 54.4
$5\times 12\times 24$ 57.1/3.4 51.2
@@
Pooling impact on Slow pathway. We analyze the effect of using different pooling strategies on both spatial and temporal dimensions over $\mathbf{F}_{v}^{\textnormal{slow}}$. The Fast pathway is kept as the default setting in Sec. 4.2. First, Table 5 (row 1 v.s. others) shows that keeping visual tokens as many as possible is a viable way to obtain better results on average, however, to cover longer context, we could easily reach the limits of an LLM’s context window and the GPU memory (e.g., the 34B model). Second, pooling properly (e.g., 2×2$\times$2 × in row 2 and 4) over either spatial or temporal dimension can also improve the performance (e.g., $\sim$1on ActivityNet-QA) but using aggressive pooling can decrease the performance a lot (row 1 v.s. row 6). This also matches the observations in prior work (Xu et al., 2024; Wu, 2024), Third, when preserving the same number of tokens (e.g., row 2 and row 4), spatial pooling obtains better results than temporal pooling, especially on the benchmarks (e.g., EgoSchema) that require strong temporal modeling capabilities.
Number of frames in Slow pathway. We experiment for the effect of using different number of frames $N_{s}$ in the Slow pathway. In particular, we test $N_{s}\in\{1,2,4,6,8,10\}$ as shown in Fig. 3, while keeping $\mathbf{F}_{v}^{\textnormal{fast}}$ in size of $50\times 4\times 4$. Note that we choose the max length to make sure the GPU memory usage of SF-LLaVA-34B inference is under 80GB. The results show that increasing the number of frames in Slow pathway can improve the performance on both ActivityNet-QA and EgoSchema. Thus we set $N^{\textnormal{slow}}$ to 10 to achieve the best possible performance of SF-LLaVA.
Pooling impact on Fast pathway. Due to the LLM’s token limit, we explore the optimal pooling strategies for the Fast pathway to model long-range context in an efficient way. We keep the Slow pathway tokens $\mathbf{F}_{v}^{\textnormal{slow}}$ in size of $10\times 12\times 24$ and $N^{\textnormal{fast}}=50$, and reduce $\mathbf{F}_{v}^{\textnormal{fast}}$ to $\{8\times 8,4\times 8,6\times 6,3\times 6,4\times 4,2\times 4,1\times 1\}$ tokens by pooling. Note that $1\times 1$ is an extreme case that loses all spatial information and can only contribute to temporal modeling. Table 6 shows that keeping more tokens, such as $8\times 8$, generally gets better results but the performance gap is not obvious. Considering that 1) SF-LLaVA-34B cannot afford more than $3\times 8$ tokens due to out-of-memory and 2) $50$ frames are able to cover videos in most datasets, we use $4\times 4$ as our default setting to trade off between spatial and temporal information. However, if we extend SF-LLaVA for understanding long-form videos over 10-20 minutes, using $1\times 1$ in the Fast pathway could be a better choice.
@@
7B $50\times 8\times 8$ 54.3/3.3 48.6
$50\times 4\times 8$ 54.5/3.3 46.8
$50\times 6\times 6$ 54.6/3.3 46.2
$50\times 3\times 8$ 54.1/3.3 46.6
$50\times 4\times 4$ 54.1/3.3 47.2
$50\times 2\times 4$ 54.3/3.3 47.4
$50\times 1\times 1$ 54.0/3.3 47.1
@@
Number of frames in Fast pathway. We evaluate the effect of using different $N^{\textnormal{fast}}$. Similar to above experiments, we keep the Slow pathway as its default in Sec. 4.2, and test to increase $N^{\textnormal{fast}}$ from 10 to 50 frames. Note that we choose 50 frames as our maximum because this reaches the GPU memory limit for the 34B model. The results in Fig. 4 show that using more frames in Fast pathway improves the performance on both ActivityNet-QA and EgoSchema datasets (e.g., using 50 frames outperforms 10 frames by 1.7% and 2.8% on EgoSchema on 7B and 34B models, respectively). Thus, SF-LLaVA uses $N^{\textnormal{fast}}=50$ in the Fast pathway.
4.5 Design Choices of Prompt
SF-LLaVA extends an Image LLM for VideoQA without further fine-tuning. Here, we explore if we should design new prompts for SF-LLaVA to better understand the new video inputs and tasks. Specifically, we decompose the prompt of Video LLMs into three main parts, “task instruction prompt”, “input data prompt”, and “structured answer prompt”. In this section, we will explain their definitions and evaluate their impact on different video tasks, ActivityNet-QA for Open-Ended VideoQA and EgoSchema for Multiple Choice VideoQA.
@@ 4.5 Design Choices of Prompt
Task instruction prompt clarifies the particular task to Video LLMs to generate desired outputs. We use ”Answer the question precisely based on the input” for Open-Ended VideoQA and Text Generation, and “Select the best option to answer the question” for Multiple Choice VideoQA. Prior work, such as LangRepo (Kahatapitiya et al., 2024), also design prompts for different types of videos (e.g., first- or third-person view), but we aim to use a prompt that can work for all videos for each category of tasks. Table 7 (row 1 and 2) shows that using the task instruction prompt can improve the performance for EgoSchema (47.2% v.s. 43.0% and 55.8%v.s. 52.2% on 7B and 34B) but is not helpful for ActivityNet-QA (55.0% v.s. 56.3% and 58.4% v.s. 59.2% on 7B and 34B). Thus, SF-LLaVA only uses the task instruction prompt for Multiple Choice VideoQA tasks.
@@
Input data prompt describes the structure of the inputs, such as the image grid in IG-VLM (Kim et al., 2024) and image sequence in PLLaVA (Xu et al., 2024). For all tasks, SF-LLaVA uses the same prompt “The input consists of a sequence of key frames from a video”. Table 7 (row 1 and 3) shows that using input data prompt obtains better results on both ActivityNet-QA (55.0% v.s. 52.6% and 58.4% v.s. 56.4%)) and EgoSchema (47.2% v.s. 44.8% and 55.8% v.s. 55.4%). This demonstrates the importance of offering input data details to better understand the structure of visual tokens, which we think is especially important to training-free methods.
@@
Structured answer prompt guides Video LLMs to generate answer in a more desirable format. This is especially important to the Multiple Choice VideoQA task, since it not only makes the answer much easier to be parsed but also improves the performance out of the box (Li et al., 2023d). In particular, we follow MVBench (Li et al., 2023d) to use “Best Option:(” as the answer prompt for Multiple Choice VideoQA and follow Image Grid (Kim et al., 2024) to use “In this video,” to guide the Open-Ended VideoQA tasks. Table 7 (row 1 and 4) shows using “structred answer prompts” improve results by 2.3% on ActivityNet-QA and 2.8% on EgoSchema with 7B LLMs.

### v2

@@ 4.3 Main Results
Open-Ended VideoQA results are shown in Table 1. SF-LLaVA obtains better performance than existing training-free methods on all benchmarks. Specifically, SF-LLaVA outperforms IG-VLM (Kim et al., 2024) by 2.1% and 5.0% on MSRVTT-QA, 5.7% and 1.5% on TGIF-QA, 1.2% and 0.8% on ActivityNet-QA, using 7B and 34B LLMs, respectively. When even compared to state-of-the-art SFT methods, SF-LLaVA achieves on-par results on most benchmarks (i.e., MSVD-QA, MSRVTT-QA, and TGIF-QA), and only the results of PLLaVA (Xu et al., 2024) and LLaVA-NeXT-Video-DPO (Zhang et al., 2024b) are better than ours on ActivityNet-QA.
Multiple Choice VideoQA results are shown in Table 2. SF-LLaVA outperforms other training-free methods that use comparable LLMs and visual encoders, such as IG-VLM (Kim et al., 2024) on all benchmarks. Specifically, on the challenging EgoSchema dataset, which involves complex long-form temporal reasoning (Mangalam et al., 2024), SF-LLaVA outperforms IG-VLM by 11.4% and 2.2% when using 7B and 34B LLMs, respectively. This highlights the ability of SF-LLaVA on long-form video understanding. Note that VideoTree (Wang et al., 2024b) is leading the benchmark because it is built upon a proprietary LLM (i.e., GPT-4 (Achiam et al., 2023)) whose performance is much better than the open-sourced LLMs. When compared to SFT methods (Cheng et al., 2024), SF-LLaVA 34B model also achieves better results (+2.5%) on EgoSchema, which confirms the capability of our SlowFast design on long videos.
@@
Text Generation benchmarks are shown in Table 3, where SF-LLaVA-34B outperforms all training-free baselines on average. First, we observe that SF-LLaVA consistently performs worse than LLaVA-NeXT-Image (Zhang et al., 2024b) on Detail Orientation (DO). This is because LLaVA-NeXT-Image takes more “high-resolution” input frames than ours (i.e., 32 frames with $12\times 12$ tokens v.s. 10 frames with $12\times 24$ tokens), thus is able to capture more spatial information. Second, SF-LLaVA takes advantage of the SlowFast design to cover a longer temporal context by using even fewer visual tokens (i.e., 4608 tokens v.s. 3680 tokens), thus excels in all other tasks, especially in Temporal Understanding (TU). Third, we observe that SF-LLaVA-34B is superior to most SFT methods (e.g, outperforming Video-LLaMA2 (Cheng et al., 2024) +0.1 score on TU and +0.31 score on CO), but only needs to catch up with LLaVA-NeXT-Video-DPO (Zhang et al., 2024b).
@@ 4.4 Design Choices of SlowFast
We first validate if both the Slow and Fast pathways are essential, and continue to experiment for their design choices respectively. These ablation studies are conducted on ActivityNet-QA (an Open-Ended VideoQA dataset that contains videos of human activities) and EgoSchema (a Multiple Choice VideoQA dataset requiring long-form understanding of egocentric videos).
Can we remove the Slow pathway? First, we simply remove the Slow pathway, while keeping the Fast pathway as 50 frames, each with $4\times 4$ tokens. Fig. 3 shows that, on all benchmarks, removing Slow pathway ($N^{\textnormal{slow}}$ equals to 0) will introduce much lower performance. Second, we validate if the performance gain is caused by the necessity of the Slow pathway or the increased visual tokens brought by using more frames. We test this by gradually increasing $N^{\textnormal{fast}}$ from 50 to 225 to compensate for the loss of visual tokens. Results in Table 4 show that using larger $N^{\textnormal{fast}}$ generally obtains better results, but the results quickly saturate when $N^{\textnormal{fast}}$ is larger than 150. We also compare the baselines in Table 4, which use $N^{\textnormal{fast}}=200$, and SF-LLaVA models in Fig. 3 with $N^{\textnormal{slow}}=8$ and $N^{\textnormal{fast}}=50$, since these models use comparable number of tokens ($3200$ v.s.3104310431043104) in total. Results show that SF-LLaVA outperforms this new baseline under all settings (e.g., 54.6% v.s. 49.7% and 46.0% v.s. 37.0% using 7B LLMs on ActivityNet-QA and EgoSchema, respectively). All of the above results demonstrate that it is essential to use the Slow pathway in SF-LLaVA.
@@
Table 4: Effect of increasing $N^{\textnormal{fast}}$ while keeping $N^{\textnormal{slow}}=0$. Each frame in the Fast pathway outputs $4\times 4$ tokens. The symbol “–” means the setting gets out-of-memory on 80GB GPUs.
Can we remove the Fast pathway? We validate this by removing the Fast pathway while retaining the Slow pathway (having 10 frames, each with $12\times 24$ tokens). Fig. 4 shows that SF-LLaVA with $N^{\textnormal{fast}}=50$ consistently outperforms this baseline. Similar to the experiments for the Slow pathway, we increase $N^{\textnormal{slow}}$ to ensure SF-LLaVA and this new baseline have a comparable number of input video tokens. Specifically, we increase $N^{\textnormal{slow}}$ to 12 frames, which is the maximum number of frames that the 34B model can afford under 80GB GPU memory. SF-LLaVA still outperforms this baseline on both ActivityNet-QA (55.5% v.s. 54.1% on 7B model and 59.2% v.s. 58.8% on 34B model) and EgoSchema (47.2% v.s. 46.6% on 7B model and 55.8% v.s. 54.6% on 34B model). We observe that the performance gap is more significant on EgoSchema, since it mostly contains long-form videos and answering the questions requires capturing longer context using the Fast pathway.
Pooling impact on Slow pathway. We analyze the effect of using different pooling strategies over $\mathbf{F}_{v}^{\textnormal{slow}}$. The Fast pathway is kept as in Sec. 4.2. First, Table 5 (row 1 v.s. others) shows that keeping visual tokens as many as possible is a viable way to obtain better results on average, however, to cover longer context, we can easily reach the limits of an LLM’s context window and the GPU memory (e.g., the 34B model). Second, pooling properly over either the spatial or temporal dimension (e.g., 2×2$\times$2 × in row 2 and 4) can also improve the performance (e.g., $\sim$1on ActivityNet-QA) but using an aggressive pooling can decrease the performance a lot (row 1 v.s. row 6). This also matches the observations in prior work (Xu et al., 2024; Wu, 2024), Third, when preserving the same number of tokens (e.g., row 2 and row 4), spatial pooling is better than temporal pooling, especially on the benchmarks (e.g., EgoSchema) that require strong temporal modeling capabilities.
Number of frames in Slow pathway. We evaluate the effect of using different numbers of frames $N_{s}$ in the Slow pathway. In particular, we test $N_{s}\in\{1,2,4,6,8,10\}$ as shown in Fig. 3, while keeping $\mathbf{F}_{v}^{\textnormal{fast}}$ in size of $50\times 4\times 4$. Note that we choose the max length to make sure the GPU memory usage of SF-LLaVA-34B inference is under 80GB. The results show that increasing the number of frames in the Slow pathway can improve the performance on both ActivityNet-QA and EgoSchema. Thus we set $N^{\textnormal{slow}}$ to 10 to achieve the best possible performance of SF-LLaVA.
Pooling impact on Fast pathway. We keep $\mathbf{F}_{v}^{\textnormal{slow}}\in\mathbb{R}^{10\times 12\times 24}$ and $N^{\textnormal{fast}}=50$, while reducing $\mathbf{F}_{v}^{\textnormal{fast}}$ to $\{8\times 8,4\times 8,6\times 6,3\times 6,4\times 4,2\times 4,1\times 1\}$ tokens by pooling. Note that $1\times 1$ is an extreme case that loses all spatial information and can only contribute to temporal modeling. Table 6 shows that keeping more tokens, such as $8\times 8$, generally gets better results on average, but the performance gap is not obvious. Considering that ($i$) SF-LLaVA-34B cannot afford more than $50\times 3\times 6$ output tokens due to out-of-memory and ($ii$) $50$ frames are able to cover most videos in existing benchmarks, we use $4\times 4$ as our default setting to trade-off between spatial and temporal information. However, if we extend SF-LLaVA for long-form video understanding (e.g., over 30 minutes), using $1\times 1$ in the Fast pathway could be a better choice to cover more input frames.
@@
7B $10\times 24\times 24$ 54.0/3.3 53.2
$10\times 12\times 24$ 55.5/3.3 47.2
$10\times 12\times 12$ 54.7/3.3 39.0
  $\;\;5\times 24\times 24$ 54.5/3.3 44.4
  $\;\;5\times 12\times 24$ 53.9/3.3 39.4
  $\;\;5\times 12\times 12$ 51.4/3.2 36.4
@@
$10\times 12\times 24$ 59.2/3.5 55.8
$10\times 12\times 12$ 58.5/3.5 50.8
  $\;\;5\times 24\times 24$ 58.3/3.5 54.4
  $\;\;5\times 12\times 24$ 57.2/3.4 51.2
@@
7B $50\times 8\times 8$ 54.9/3.3 48.6
$50\times 4\times 8$ 55.0/3.3 46.8
$50\times 6\times 6$ 55.2/3.3 46.2
$50\times 3\times 6$ 54.9/3.3 46.6
$50\times 4\times 4$ 55.5/3.3 47.2
$50\times 2\times 4$ 54.7/3.3 47.4
$50\times 1\times 1$ 54.4/3.3 47.1
@@
Number of frames in Fast pathway. We evaluate the effect of using different $N^{\textnormal{fast}}$. Similar to the above experiments, we keep the Slow pathway as its default in Sec. 4.2, and test to increase $N^{\textnormal{fast}}$ from 10 to 50 frames. Note that we chose 50 frames as our maximum because this reaches the GPU memory limit for the 34B model. The results in Fig. 4 show that using more frames in Fast pathway improves the performance on both ActivityNet-QA and EgoSchema datasets (e.g., using 50 frames outperforms 10 frames by 1.7% and 2.8% on EgoSchema on 7B and 34B models, respectively). Thus, by default, SF-LLaVA uses $N^{\textnormal{fast}}=50$ in the Fast pathway.
4.5 Design Choices of Prompt
SF-LLaVA is built upon a pre-trained Image LLM for VideoQA without further fine-tuning. Here we evaluate if we should design new prompts for SF-LLaVA to better understand the video inputs and tasks. We decompose the prompt into three main parts, and respectively explore their best designs.
@@ 4.5 Design Choices of Prompt
Task instruction prompt clarifies the goal of the target task. In particular, we use “Answer the question precisely based on the input” for the open-ended task, and “Select the best option to answer the question” for the multiple choice task. Table 7 (row 1 and 2) shows that using the task instruction prompt can improve the performance on EgoSchema (47.2% v.s. 43.0% and 55.8%v.s. 52.2% on 7B and 34B) but is not helpful for ActivityNet-QA (54.9% v.s. 55.5% and 58.4% v.s. 59.2% on 7B and 34B). Thus, SF-LLaVA uses the task instruction prompt only for Multiple Choice VideoQA.
@@
Input data prompt describes the structure of the inputs, such as the image grid in IG-VLM (Kim et al., 2024) and image sequence in PLLaVA (Xu et al., 2024). For all tasks, SF-LLaVA uses the same prompt “The input consists of a sequence of key frames from a video”. Table 7 (row 1 and 3) shows that using input data prompt obtains better results on both ActivityNet-QA (54.9% v.s. 52.6% and 58.4% v.s. 56.4%) and EgoSchema (47.2% v.s. 44.8% and 55.8% v.s. 55.4%). This demonstrates the importance of offering input data details to better understand the structure of visual tokens, which we think is especially important to training-free methods.
@@
Structured answer prompt guides Video LLMs to generate answers in a more desirable format. This is especially important to the Multiple Choice VideoQA task, since it makes the answer easier to be parsed and improves the performance out of the box (Li et al., 2023d). We follow MVBench (Li et al., 2023d) to use “Best Option:(” as the answer prompt for Multiple Choice VideoQA and follow Image Grid (Kim et al., 2024) to use “In this video,” to guide the Open-Ended VideoQA tasks. Table 7 (row 1 and 4) shows that using structured answer prompts improves results by 2.2% on ActivityNet-QA and 2.8% on EgoSchema with the 7B LLM.
Figure 5: Qualitative results of SF-LLaVA-7B. Our model can correctly understand the video content and capture important details. On the other hand, error cases show that SF-LLaVA lacks the capability of fine-grained temporal grounding and may miss certain moments due to uniform sampling.
4.6 Error Analysis
First, although SF-LLaVA can understand the relative sequence of different events, it still lacks the capability to detect the precise start and end time of a video moment, such as the Question $2$ in Fig. 5. This is because SF-LLaVA is never trained on any video datasets and relevant tasks. Fine-tuning SF-LLaVA on fine-grained time-related video datasets could be a promising direction to gain this capability, and incorporating multimodal inputs (e.g., the timestamp, subtitle, and audio of each frame) can further improve the performance. Second, as observed in many other methods, SF-LLaVA possibly misses some key frames due to its uniform frame sampling. Question $3$ in Fig. 5 is an example, in which the frames showing a quick moment of “opening refrigerator” are unintentionally missed (thus not shown in Fig. 5). Drastically sampling more frames can mitigate this issue but is limited by an LLM’s context window. This suggests we may explore dynamic sampling strategies to ensure a more comprehensive sampling of important video segments.